OpenStudy (anonymous):

Search Results What is the simplified form of 2x + 4 / x -2 minus x - 3/ x-2 ? is it x+1/x-2

OpenStudy (unklerhaukus):

almost \[\frac{2x + 4 }{ x -2}- \frac{x - 3}{ x-2}=\frac{2x+4-(x-3)}{x-2}=\]

OpenStudy (anonymous):

x+7/x-2 ?

OpenStudy (anonymous):

\[\frac{ x+7 }{ x-2 }\] yes thats right

OpenStudy (unklerhaukus):

thats better

OpenStudy (anonymous):

thankyou guys . if i need any more help. can i tagg you ?

OpenStudy (anonymous):

of course

OpenStudy (unklerhaukus):

you can but i wont be online much more today/tonight, you can simplify further

OpenStudy (unklerhaukus):

\[\frac{ x+7 }{ x-2 }=\frac{ x-2+9 }{ x-2 }=\]

OpenStudy (anonymous):

x+7/x-2 ?

OpenStudy (anonymous):

the same thhing right ?

OpenStudy (unklerhaukus):

\[\frac{ x+7 }{ x-2 }=\frac{ x-2+9 }{ x-2 }=\frac{x-2}{x-2}+\frac{9}{x-2}=\]

OpenStudy (unklerhaukus):

...

OpenStudy (anonymous):

i dont knoe :/

OpenStudy (unklerhaukus):

you can simplify the \(\frac{x-2}{x-2}\) bit